OpenStudy (anonymous):

Find the monthly payment for the loan. (Round your answer to the nearest cent.) $100 loan for 12 months at 15%

OpenStudy (tkhunny):

\(i = 0.15\) \(j = I/12\) \(v = 1/(1+j)\) \(Pmt = 100\cdot\dfrac{1-v}{v-v^{13}}\) If you need help with that, please supply formulas and some discussion from your course materials. Generally, don't post a problem with no work shown.